OpenStudy (anonymous):

WILL STAR & FAN! \[\sqrt{p^2-3p+49}=p+4\] the answer is 3 but I don't get how it was gotten I tried squaring both sides and then canceling out and the answer i got was 11...

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

What happened when you squared both sides?

OpenStudy (anonymous):

\[ p^2-3p+49 = p^2+8p+16 \]

OpenStudy (anonymous):

\[ -3p+49 =8p+16 \]

OpenStudy (anonymous):

\[ 49 =11p+16 \]

OpenStudy (anonymous):

\[ 33 =11p \]

OpenStudy (anonymous):

\[ p=3 \]

OpenStudy (anonymous):

@Procrasturbator